Twenty years​ ago, 53​% of parents of children in high school felt it was a serious problem that high school students were not being taught enough math and science. A recent survey found that 310 of 850 parents of children in high school felt it was a serious problem that high school students were not being taught enough math and science. Do parents feel differently today than they did twenty years​ ago? Use the alpha equals α=0.05 level of significance.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.53

Ha : p 0.53

= x / n = 310 /850 = 0.3647

P0 = 0.53

1 - P0 = 1 -0.53 = 0.47

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.3647 - 0.53 / [0.53*(0.47) /850 ]

= -9.656

P(z < -9.656) = 0

P-value = 0

= 0.05    

0 < 0.05

Reject the null hypothesis .

There is sufficient evidence to suggest that   

Yes.Do parents feel differently today than they did twenty years​ ago.